# A priori error analysis of the proximal Galerkin method

### v1

@@ Abstract.
The proximal Galerkin (PG) method is a finite element method for solving variational problems with inequality constraints. It has several advantages, including constraint-preserving approximations and mesh independence. This paper presents the first abstract a priori error analysis of PG methods, providing a general framework to establish convergence and error estimates. As applications of the framework, we demonstrate optimal convergence rates for both the obstacle and Signorini problems using various finite element subspaces.
@@ 1.2. Notation
[…]
@@
where $\tilde{w}$ is the extension by zero of $w$ outside of $\Gamma$; i.e., $\tilde{w}=0$ on $\partial\Omega\backslash\overline{\Gamma}$ and $\tilde{w}=w$ on $\Gamma$. We note that this space is normed by $\|w\|_{\widetilde{H}^{1/2}(\Gamma)}:=\|\tilde{w}\|_{H^{1/2}(\partial\Omega)}$ and that ${H}^{-1/2}(\Gamma)=(\widetilde{H}^{1/2}(\Gamma))^{\prime}$. For non-integer $s$, the notation $H^{s}(\Omega)$ denotes the Sobolev–Slobodeckij spaces [24, Chapter 2]. We use the notation $(\cdot,\cdot)_{\omega}$ to denote the $L^{2}(\omega)$-inner product over measurable $\omega\subset\overline{\Omega}$. If $\omega=\Omega$, we drop the subscript and denote the $L^{2}$-inner product over $\Omega$ by (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ).
@@ 1.3. General setup
for positive constants $\nu$ and $M$, and $F$ is a bounded linear functional on $V$. The convexity of $K$ implies that the model problem 4 is equivalent to the following variational inequality [18, Theorem 6.1-2]: find $u^{*}\in K$, such that
@@
Owing to the coercivity of $a$, 9 admits a unique solution [11, Theorem 5.6]. Moreover, $E(u^{\ast})\leq E(v)$ for all $v\in K$.
@@ 1.4. The proximal Galerkin method
[…]
@@ 1.5. Main results
  • Finally, one of the main results of this paper may be summarized by the following estimate. If $V\subset H^{1}(\Omega)$ and the solution $u^{*}\in H^{1+s}(\Omega)$ with $E^{\prime}(u^{*})\in H^{1-r}(\Omega)$ for some $s,r\in(0,1]$, then
    (15) $\|u^{*}-u_{h}^{\ell}\|^{2}_{H^{1}(\Omega)}+\|\lambda^{*}-\lambda_{h}^{\ell}\|^{2}_{Q^{\prime}}\lesssim\frac{C_{\rm{stab}}}{\sum_{k=1}^{\ell}\alpha_{k}}+C_{\mathrm{reg}}h^{2\cdot\min\{r,s\}}$
@@ 1.6. Example problems
Korn’s inequality guarantees the coercivity of the bilinear form $a$ over $V$; see, e.g., [26, Theorem 42.9]. In this case, $\lambda^{*}=(\mathsf{C}\epsilon(u^{*})n)\cdot n\in{H}^{-1/2}(\Gamma_{\mathrm{T}})=Q^{\prime}$.
@@
Note that $a$ is coercive over the Hilbert space $V$ due to a Friedrichs’ inequality; see, e.g., [20, Lemma 2.8].
@@ 1.7. Closed observation maps: A conjecture
Note that this set is recovered from 5 by setting $C$ to be the closed unit ball in $\mathbb{R}^{n}$, $V=H^{1}_{0}(\Omega)$, $B=\nabla$, and $\Omega_{d}=\Omega$. For further details, see [21, Example 6]. If $n\geq 2$, then $\operatorname{im}B$ is a closed proper subspace of $L^{2}(\Omega;\mathbb{R}^{n})$.
@@ 2.1. Legendre functions
In turn, we conclude that $\operatorname{dom}(\nabla\mathcal{R})=\operatorname{im}(\nabla\mathcal{R}^{*})\subset\mathcal{O}$, implying $\nabla\mathcal{R}^{*}(\psi)(x)\in\operatorname{int}C(x)$ f.a.e. $x\in\Omega_{d}$. We provide three explicit examples in Examples 2.1, 2.2 and 2.3 below, and we refer to [21, Table 1] for more.
@@
This entropy has been used in the PG framework for gradient norm constraints, see Section 1.7 and [21, Example 6].
@@ 2.2. Bregman divergences
The proof can be found in [2], see also [6, Theorem 3.9]. We also recall the three points identity [15, Lemma 3.1]:
@@ 2.3. Deriving the proximal Galerkin method
(38) find $\text{find }u^{k}\in K\text{ such that }~{}\alpha_{k}\langle E^{\prime}(u^{k}),v\rangle+(\nabla\mathcal{R}(Bu^{k}),Bv)_{\Omega_{d}}=(\nabla\mathcal{R}(Bu^{k-1}),Bv)_{\Omega_{d}}$
@@
[…]
@@ 3.1. Compatible subspaces
see, e.g., [26, Lemma 26.9] and [8, Section 5.4.3]. In what follows, we let $\|\Pi_{h}\|$ denote the operator norm of $\Pi_{h}$.
@@
The following inf-sup condition holds with a constant $\beta>0$ [16, Proposition 7.2 and Remark 7.2]:
@@
with $Q=L^{2}(\Omega)$. For the Legendre function $\mathcal{R}$, one can choose the Fermi–Dirac entropy given in Example 2.2 with $\underline{u}=-1$ and $\overline{u}=1$, although other convenient choices are also appropriate; cf. [28]. The continuous LBB condition 41 holds thanks to the surjectivity of the divergence operator from $H(\mathrm{div},\Omega)$ to $L^{2}(\Omega)$; see, e.g., [26, Lemma 51.2]. To ensure that 40 holds, a natural choice of subspace would be the $H(\mathrm{div})$-conforming Raviart–Thomas space for $V_{h}$ and the broken polynomial space of the same order for $W_{h}$; see, e.g., [26, Lemma 51.10].
@@ 3.2. The PG method is well-defined
where, in the first step, we used that the isomorphism $A^{-1}:V_{h}^{\prime}\to V_{h}$ is injective with closed range. In the last estimate, we used 48. Together with the fact that we can lower bound $(\mathcal{R}^{*}(w),1)_{\Omega_{d}}$ by an affine linear function and the equivalence of norms in finite dimensional spaces, we deduce that $-J(w)\to\infty$ as $\|w\|_{Q^{\prime}}\to\infty$. The map $w\mapsto v(w)$ is continuous as the first term is affine linear and $\mathcal{R}^{*}$ is continuous, hence we can guarantee the existence of a maximizer $w^{*}$ for $J$.
We now show that the pair $(v(w^{*}),w^{*})$ is the sought-after saddle point of $\mathcal{L}$. This follows from standard arguments; see e.g., [18, Exercise 7.16-4]. We provide some details for completeness. First observe that since $\mathcal{L}(v,\cdot)$ is concave, we have that
@@ 3.3. Energy dissipation
Observe that from 10b,
(57) $b(u_{h}^{k+1}-u_{h}^{k},w_{h})-(\nabla\mathcal{R}^{*}(\psi_{h}^{k+1})-\nabla\mathcal{R}^{*}(\psi_{h}^{k}),w_{h})_{\Omega_{d}}=0~{}\text{ for all }w_{h}\in W_{h}.$
Since $E$ is convex, we obtain that for all $k\geq 1$
$E(u_{h}^{k+1})\leq E(u_{h}^{k})+\langle E^{\prime}(u_{h}^{k+1}),u_{h}^{k+1}-u_{h}^{k}\rangle.$
Using the definition of $E^{\prime}(u)\in V^{\prime}$:
(58) $\langle E^{\prime}(u),v\rangle=a(u,v)-F(v)~{}\text{ for all }u,v\in V,$
along with 10a and 57 yields
(59) $E(u_{h}^{k+1})$ $\leq E(u_{h}^{k})+\frac{1}{\alpha_{k+1}}b(u_{h}^{k+1}-u_{h}^{k},\psi_{h}^{k}-\psi_{h}^{k+1})$
@@
With 36, we obtain the result. ∎
@@ 3.4. Best approximation error estimates
The error estimates of this section comprise an optimization error, governed by the sequence $\{\alpha_{k}\}$, and an discretization error, determined by the discretization choice. The analysis requires constructing a specific enriching map and a nonlinear operator defined in 65, below. In turn, we make the following general assumption, which is critical to defining the operator. First, we motivate this assumption with a remark.
Remark 1 (Motivating Assumption 2).
As expected from the analysis of nonconforming discretizations of obstacle problems [5, Section 5.2.1], one then requires an operator $\mathcal{E}_{h}:K_{h}\rightarrow K$ with suitable approximation properties. For the analysis of PG, we generally require that $\mathcal{E}_{h}$ maps to $\operatorname{dom}(\nabla\mathcal{R}\circ B)\subset K$. This allows us to define another approximation $\mathcal{U}_{h}$ that is compatible with the nonlinear term in 10b. See Section 4.3 for an explicit construction of the enriching map for $K_{h}$ defined in 60.
Assumption 2.
(61) $\mathcal{E}_{h}:K\cup K_{h}\rightarrow\operatorname{dom}(\nabla\mathcal{R}\circ B)\subset K,$
where
The set $K_{h}$ contains functions $u_{h}\in V_{h}$ such that $Bu_{h}\in W_{h}$ is the $L^{2}$-projection of a constrained observable $o\in\mathcal{O}$. It is constructed to ensure that $u_{h}^{k}\in K_{h}$ for all iterations $k$; see 10b. Under Assumption 2, we associate to any $u\in K\cup K_{h}$ a function
(64) $\psi_{u}:=(\nabla\mathcal{R}\circ B)(\mathcal{E}_{h}u).$
Finally, we denote by
$\Psi_{h}=\{\psi\in L^{\infty}(\Omega_{d};\mathbb{R}^{m})\mid\psi=\psi_{u}\text{ for some }u\in K\cup K_{h}\},$
and define the nonlinear operator $\mathcal{U}_{h}:\Psi_{h}\rightarrow V_{h}$
(65) $\mathcal{U}_{h}(\psi_{u})=\Pi_{h}(\mathcal{E}_{h}u).$
Observe that by the definition of the Fortin operator 42 and the identity $\nabla\mathcal{R}^{*}=(\nabla\mathcal{R})^{-1}$ in 26, it follows that
(66) $b(\mathcal{U}_{h}(\psi_{u}),w_{h})=b(\mathcal{E}_{h}u,w_{h})=(\nabla\mathcal{R}^{*}(\psi_{u}),w_{h})_{\Omega_{d}}\quad~{}\text{ for all }w_{h}\in W_{h}.$
Furthermore, the nonlinear operator $\mathcal{U}_{h}$ allows us to derive a key identity stated in Lemma 3.3.
Lemma 3.3 (Identity).
Let Assumption 2 hold and let $\psi\in\Psi_{h}$ be given. For any $k\geq 1$, we have that
(67) $E(u_{h}^{k})-E(\mathcal{U}_{h}(\psi))$ $\leq\langle E^{\prime}(u_{h}^{k}),u_{h}^{k}-\mathcal{U}_{h}(\psi)\rangle\leq\frac{1}{\alpha_{k}}(\mathcal{D}^{*}(\psi_{h}^{k-1},\psi)-\mathcal{D}^{*}(\psi_{h}^{k},\psi),1)_{\Omega_{d}}.$
@@
For any $k\geq 1$, we use 10a with $v_{h}=u_{h}^{k}-\mathcal{U}_{h}(\psi)$ and write
$\langle E^{\prime}(u_{h}^{k}),u_{h}^{k}-\mathcal{U}_{h}(\psi)\rangle=\frac{1}{\alpha_{k}}b(u_{h}^{k}-\mathcal{U}_{h}(\psi),\psi_{h}^{k-1}-\psi_{h}^{k}).$
We use 10b, property 66, and the three points identity 36 to proceed:
$b(u_{h}^{k}-\mathcal{U}_{h}(\psi),\psi_{h}^{k-1}-\psi_{h}^{k})$ $=(\psi_{h}^{k-1}-\psi_{h}^{k},\nabla\mathcal{R}^{*}(\psi_{h}^{k})-\nabla\mathcal{R}^{*}(\psi))_{\Omega_{d}}$
$=(-\mathcal{D}^{*}(\psi_{h}^{k},\psi)+\mathcal{D}^{*}(\psi_{h}^{k-1},\psi)-\mathcal{D}^{*}(\psi_{h}^{k-1},\psi_{h}^{k}),1)_{\Omega_{d}}$
$\leq(-\mathcal{D}^{*}(\psi_{h}^{k},\psi)+\mathcal{D}^{*}(\psi_{h}^{k-1},\psi),1)_{\Omega_{d}},$
where, in the last line, we have used that $\mathcal{D}^{*}(\psi_{h}^{k-1},\psi_{h}^{k})\geq 0$. In summary, we conclude that
(68) $\langle E^{\prime}(u_{h}^{k}),u_{h}^{k}-\mathcal{U}_{h}(\psi)\rangle\leq\frac{1}{\alpha_{k}}(\mathcal{D}^{*}(\psi_{h}^{k-1},\psi)-\mathcal{D}^{*}(\psi_{h}^{k},\psi),1)_{\Omega_{d}}.$
The convexity of $E$ gives
$E(u_{h}^{k})-E(\mathcal{U}_{h}(\psi))\leq\langle E^{\prime}(u_{h}^{k}),u_{h}^{k}-\mathcal{U}_{h}(\psi)\rangle.$
Applying 68 to the above shows 67. ∎
We are now ready to derive the main best approximation result.
Theorem 3.4 (Best approximation of the primal iterates $u_{h}^{\ell}$).
Let $u^{*}\in K$ be the solution to 4, let $\{u_{h}^{k}\}_{k=1}^{\ell}$ be defined via Algorithm 1. Assume we are in the setting outlined in Sections 1.3 and 3.1 and let Assumption 2 hold. Then the following estimate is valid for every $\ell\geq 1$:
(69) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}\\ \leq\inf_{\psi\in\Psi_{h},v\in K}\left(\frac{(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|u^{\ast}-\mathcal{U}_{h}(\psi)\|_{V}^{2}+|\langle E^{\prime}(u^{*}),u_{h}^{\ell}-v+u^{*}-\mathcal{U}_{h}(\psi)\rangle|\right).$
Proof.
Recall the definition of $E^{\prime}$ 58 and observe that for any $u,w\in V$
(70) $E(u)-E(w)\geq\langle E^{\prime}(w),u-w\rangle+\frac{1}{2}a(u-w,u-w)\geq\langle E^{\prime}(w),u-w\rangle+\frac{\nu}{2}\|u-w\|^{2}_{V},$
where we used the coercivity property given in 7. For any $\psi\in\Psi_{h}$ and $u\in V$, we arrive at the following inequality when setting $w=u_{h}^{k}$ above:
@@
$=E(u_{h}^{k})+\langle E^{\prime}(u_{h}^{k}),\mathcal{U}_{h}(\psi)-u_{h}^{k}\rangle+\langle E^{\prime}(u_{h}^{k}),u-\mathcal{U}_{h}(\psi)\rangle+\frac{\nu}{2}\|u-u_{h}^{k}\|_{V}^{2}.$
Applying Lemma 3.3 to the second term above yields
(71) $E(u)\geq E(u_{h}^{k})+\frac{1}{\alpha_{k}}(\mathcal{D}^{*}(\psi_{h}^{k},\psi)-\mathcal{D}^{*}(\psi_{h}^{k-1},\psi),1)_{\Omega_{d}}\\ +\langle E^{\prime}(u_{h}^{k}),u-\mathcal{U}_{h}(\psi)\rangle+\frac{\nu}{2}\|u-u_{h}^{k}\|_{V}^{2}.$
@@
We now multiply 71 by $\alpha_{k}$, sum from $k=1$ to $k=\ell$, use the energy dissipation property Lemma 3.2, and rearrange the resulting inequality. We obtain
(72) $(E(u_{h}^{\ell})-E(u))\sum_{k=1}^{\ell}\alpha_{k}+\sum_{k=1}^{\ell}\frac{\alpha_{k}}{2}\nu\|u-u_{h}^{k}\|_{V}^{2}+(\mathcal{D}^{*}(\psi_{h}^{\ell},\psi),1)_{\Omega_{d}}\\ \leq(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}+\sum_{k=1}^{\ell}\alpha_{k}\langle E^{\prime}(u_{h}^{k}),\mathcal{U}_{h}(\psi)-u\rangle.$
@@
Then, upon dividing 72 by $\sum_{k=1}^{\ell}\alpha_{k}$, followed by adding a subtracting $\langle E^{\prime}(u),\mathcal{U}_{h}(\psi)-u\rangle$ from the right-hand side of 72, we find that
$E(u_{h}^{\ell})-E(u)+\frac{1}{2}\frac{\sum_{k=1}^{\ell}\alpha_{k}\nu\|u-u_{h}^{k}\|_{V}^{2}}{\sum_{k=1}^{\ell}\alpha_{k}}\\ \leq\frac{(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\langle E^{\prime}(u),\mathcal{U}_{h}(\psi)-u\rangle+\frac{\sum_{k=1}^{\ell}\alpha_{k}\langle E^{\prime}(u_{h}^{k})-E^{\prime}(u),\mathcal{U}_{h}(\psi)-u\rangle}{\sum_{k=1}^{\ell}\alpha_{k}}.$
We now define the weighted average $\overline{u}_{h}^{\ell}=\sum_{k=1}^{\ell}\alpha_{k}u_{h}^{k}/\sum_{k=1}^{\ell}\alpha_{k}$. Using Jensen’s inequality, we conclude that
(73) $\|u-\overline{u}_{h}^{\ell}\|_{V}^{2}\leq\frac{\sum_{k=1}^{\ell}\alpha_{k}\|u-u_{h}^{k}\|_{V}^{2}}{\sum_{k=1}^{\ell}\alpha_{k}}$
and, in turn,
$E(u_{h}^{\ell})$ $-E(u)+\frac{\nu}{2}\|u-\overline{u}_{h}^{\ell}\|_{V}^{2}$

### v2

@@ Abstract.
The proximal Galerkin (PG) method is a finite element method for solving variational problems with inequality constraints. It has several advantages, including constraint-preserving approximations and mesh independence. This paper presents the first abstract a priori error analysis of the PG method, providing a general framework to establish convergence and error estimates. As applications of the framework, we demonstrate optimal convergence rates for both the obstacle and Signorini problems using various finite element subspaces.
@@ 1.2. Notation
Throughout the article, we let $\Omega\subset\mathbb{R}^{n}$ ($n=1,2,3$) be an open bounded Lipschitz domain. For a given Banach space $V$, we denote by $V^{\prime}$ its topological dual space with duality pairing $\langle\cdot,\cdot\rangle$. In particular, a member $F\in V^{\prime}$ is a continuous linear functional mapping $V$ into $\mathbb{R}$, $F(v)=\langle F,v\rangle\in\mathbb{R}$. The norm $\|\cdot\|_{V^{\prime}}$ denotes the usual operator norm.
@@
where $\tilde{w}$ is the extension by zero of $w$ outside of $\Gamma$; i.e., $\tilde{w}=0$ on $\partial\Omega\backslash\overline{\Gamma}$ and $\tilde{w}=w$ on $\Gamma$. We note that this space is normed by $\|w\|_{\widetilde{H}^{1/2}(\Gamma)}:=\|\tilde{w}\|_{H^{1/2}(\partial\Omega)}$ and that ${H}^{-1/2}(\Gamma)=(\widetilde{H}^{1/2}(\Gamma))^{\prime}$. For non-integer $s$, the notation $H^{s}(\Omega)$ denotes the Sobolev–Slobodeckij spaces [23, Chapter 2]. We use the notation $(\cdot,\cdot)_{\omega}$ to denote the $L^{2}(\omega)$-inner product over measurable $\omega\subset\overline{\Omega}$. If $\omega=\Omega$, we drop the subscript and denote the $L^{2}$-inner product over $\Omega$ by $(\cdot,\cdot)$.
@@ 1.3. General setup
for positive constants $\nu$ and $M$, and $F$ is a bounded linear functional on $V$. The convexity of $K$ implies that the model problem 4 is equivalent to the following variational inequality [16, Theorem 6.1-2]: find $u^{*}\in K$, such that
@@
Owing to the coercivity of $a$, 9 admits a unique solution [10, Theorem 5.6]. Moreover, $E(u^{\ast})\leq E(v)$ for all $v\in K$.
@@ 1.4. The proximal Galerkin method
(10a) $\alpha_{k}\,a(u^{k}_{h},v_{h})+b(v_{h},\psi^{k}_{h}-\psi_{h}^{k-1})$ $=\alpha_{k}\,F(v_{h})$  $~\text{ for all }v_{h}\in V_{h},$
(10b) $b(u^{k}_{h},w_{h})-(\nabla\mathcal{R}^{*}(\psi^{k}_{h}),w_{h})_{\Omega_{d}}$ $=0$  $~\text{ for all }w_{h}\in W_{h}.$
@@ 1.5. Main results
  • Finally, one of the main results of this paper may be summarized by the following estimate. If $V\subset H^{1}(\Omega)$ and the solution $u^{*}\in H^{1+s}(\Omega)$ with $E^{\prime}(u^{*})\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H^{r-1}(\Omega)}$ for some $s,r\in(0,1]$, then
    (15) $\|u^{*}-u_{h}^{\ell}\|^{2}_{H^{1}(\Omega)}+\|\lambda^{*}-\lambda_{h}^{\ell}\|^{2}_{Q^{\prime}}\leq\frac{C_{\rm{stab}}}{\sum_{k=1}^{\ell}\alpha_{k}}+C_{\mathrm{reg}}h^{2\cdot\min\{r,s\}}$
@@ 1.6. Example problems
Korn’s inequality guarantees the coercivity of the bilinear form $a$ over $V$; see, e.g., [25, Theorem 42.9]. In this case, $\lambda^{*}=(\mathsf{C}\epsilon(u^{*})n)\cdot n\in{H}^{-1/2}(\Gamma_{\mathrm{T}})=Q^{\prime}$.
@@
Note that $a$ is coercive over the Hilbert space $V$ due to a Friedrichs’ inequality; see, e.g., [18, Lemma 2.8].
@@ 1.7. Closed observation maps: A conjecture
Note that this set is recovered from 5 by setting $C$ to be the closed unit ball in $\mathbb{R}^{n}$, $V=H^{1}_{0}(\Omega)$, $B=\nabla$, and $\Omega_{d}=\Omega$. For further details, see [19, Example 6]. If $n\geq 2$, then $\operatorname{im}B$ is a closed proper subspace of $L^{2}(\Omega;\mathbb{R}^{n})$.
@@ 2.1. Legendre functions
In turn, we conclude that $\operatorname{dom}(\nabla\mathcal{R})=\operatorname{im}(\nabla\mathcal{R}^{*})\subset\mathcal{O}$, implying $\nabla\mathcal{R}^{*}(\psi)(x)\in\operatorname{int}C(x)$ f.a.e. $x\in\Omega_{d}$. We provide three explicit examples in Examples 2.1, 2.2 and 2.3 below, and we refer to [19, Table 1] for more.
@@
This entropy has been used in the PG framework for gradient norm constraints, see Section 1.7 and [19, Example 6].
@@ 2.2. Bregman divergences
The proof can be found in [2], see also [6, Theorem 3.9]. We also recall the three points identity [13, Lemma 3.1]:
@@ 2.3. Deriving the proximal Galerkin method
(38) $\text{find }u^{k}\in K\text{ such that }~\alpha_{k}\langle E^{\prime}(u^{k}),v\rangle+(\nabla\mathcal{R}(Bu^{k}),Bv)_{\Omega_{d}}=(\nabla\mathcal{R}(Bu^{k-1}),Bv)_{\Omega_{d}}$
@@
(39a) $\alpha_{k}a(u^{k},v)+b(v,\psi^{k}-\psi^{k-1})$ $=\alpha_{k}F(v)$  $~\text{ for all }v\in V,$
(39b) $b(u^{k},w)-(\nabla\mathcal{R}^{*}(\psi^{k}),w)_{\Omega_{d}}$ $=0$  $~\text{ for all }w\in W,$
@@ 3.1. Compatible subspaces
see, e.g., [25, Lemma 26.9] and [7, Section 5.4.3]. In what follows, we let $\|\Pi_{h}\|$ denote the operator norm of $\Pi_{h}$.
@@
The following inf-sup condition holds with a constant $\beta>0$ [14, Proposition 7.2 and Remark 7.2]:
@@
with $Q=L^{2}(\Omega)$. For the Legendre function $\mathcal{R}$, one can choose the Fermi–Dirac entropy given in Example 2.2 with $\underline{u}=-1$ and $\overline{u}=1$, although other convenient choices are also appropriate; cf. [28]. The continuous LBB condition 41 holds thanks to the surjectivity of the divergence operator from $H(\mathrm{div},\Omega)$ to $L^{2}(\Omega)$; see, e.g., [25, Lemma 51.2]. To ensure that 40 holds, a natural choice of subspace would be the $H(\mathrm{div})$-conforming Raviart–Thomas space for $V_{h}$ and the broken polynomial space of the same order for $W_{h}$; see, e.g., [25, Lemma 51.10].
@@ 3.2. The PG method is well-defined
where, in the first step, we used that the isomorphism $A^{-1}:V_{h}^{\prime}\to V_{h}$ is injective with closed range. In the last estimate, we used 48. Together with the fact that we can lower bound $(\mathcal{R}^{*}(w),1)_{\Omega_{d}}$ by an affine linear function and the equivalence of norms in finite dimensional spaces, we deduce that $J(w)\to\infty$ as $\|w\|_{Q^{\prime}}\to\infty$. The map $w\mapsto v(w)$ is continuous as the first term is affine linear and $\mathcal{R}^{*}$ is continuous, hence we can guarantee the existence of a maximizer $w^{*}$ for $J$.
We now show that the pair $(v(w^{*}),w^{*})$ is the sought-after saddle point of $\mathcal{L}$. This follows from standard arguments; see e.g., [16, Exercise 7.16-4]. We provide some details for completeness. First observe that since $\mathcal{L}(v,\cdot)$ is concave, we have that
@@ 3.3. Energy dissipation
Utilizing the coercivity of $a$ (8) and (10a)–(10b), we directly observe that
(57) $E(u_{h}^{k+1})$ $\leq E(u_{h}^{k})+\langle E^{\prime}(u_{h}^{k+1}),u_{h}^{k+1}-u_{h}^{k}\rangle$
$=E(u_{h}^{k})+\frac{1}{\alpha_{k+1}}b(u_{h}^{k+1}-u_{h}^{k},\psi_{h}^{k}-\psi_{h}^{k+1})$
@@
The three points identity (36) provides the result. ∎
@@ 3.4. Best approximation error estimates
Theorem 3.3 (Best approximation of the primal iterates $u_{h}^{\ell}$).
Let $u^{*}\in K$ be the solution to 4, let $\{u_{h}^{k}\}_{k=1}^{\ell}$ be defined via Algorithm 1. Assume we are in the setting outlined in Sections 1.3 and 3.1. Then the following estimate is valid for every $\ell\geq 1$:
(58) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}\leq\frac{(\mathcal{D}(o^{*},o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|\Pi_{h}u^{*}-u^{*}\|^{2}_{V}\\ +|\langle E^{\prime}(u^{*}),\Pi_{h}u^{*}-u^{*}\rangle|+\inf_{v\in K}|\langle E^{\prime}(u^{*}),v-u_{h}^{\ell}\rangle|,$
where we recall that $o^{*}=Bu^{*}$, $o_{h}^{0}=\nabla\mathcal{R}^{*}(\psi_{h}^{0})$, and $\Pi_{h}$ is the Fortin map given in (42).
@@
Observe that with the coercivity property of $a$ (8), we have for any $u,v\in V$
(59) $E(u)-E(v)=\langle E^{\prime}(v),u-v\rangle+\frac{1}{2}a(u-v,u-v)\geq\langle E^{\prime}(u),u-v\rangle+\frac{\nu}{2}\|u-v\|^{2}_{V},$
Setting $v=u_{h}^{k}$ above, we arrive at the following inequality for any $u\in V$:
@@
$=E(u_{h}^{k})+\langle E^{\prime}(u_{h}^{k}),\Pi_{h}u-u_{h}^{k}\rangle+\langle E^{\prime}(u_{h}^{k}),u-\Pi_{h}u\rangle+\frac{\nu}{2}\|u-u_{h}^{k}\|_{V}^{2}.$
We now rewrite the second term above using (10a), the definition of the dual variable $\lambda_{h}^{k}$ (13), and the Fortin map (42). We obtain
(60) $\langle E^{\prime}(u_{h}^{k}),\Pi_{h}u-u_{h}^{k}\rangle=b(\Pi_{h}u-u_{h}^{k},\lambda_{h}^{k})=b(u-u_{h}^{k},\lambda_{h}^{k}).$
Proceeding, we use some ideas from [21, 27]. Namely, we use (10b), the definition of $o_{h}^{k}$ (12), the observation that $\psi_{h}^{k}=\nabla\mathcal{R}(o_{h}^{k})$ from (26), and the three points identity (36). This yields
(61) $b(u-u_{h}^{k},\lambda_{h}^{k})=(Bu-Bu_{h}^{k},\lambda_{h}^{k})_{\Omega_{d}}$ $=\frac{1}{\alpha_{k}}(Bu-o_{h}^{k},\nabla\mathcal{R}(o_{h}^{k-1})-\nabla\mathcal{R}(o_{h}^{k}))_{\Omega_{d}}$
$=\frac{1}{\alpha_{k}}(\mathcal{D}(Bu,o_{h}^{k})-\mathcal{D}(Bu,o_{h}^{k-1})+\mathcal{D}(o_{h}^{k},o_{h}^{k-1}),1)_{\Omega_{d}}.$
Collecting the above and using $\mathcal{D}(o_{h}^{k},o_{h}^{k-1})\geq 0$, we obtain the following identity:
(62) $E(u_{h}^{k})-E(u)+\frac{1}{\alpha_{k}}(\mathcal{D}(Bu,o_{h}^{k})-\mathcal{D}(Bu,o_{h}^{k-1}),1)_{\Omega_{d}}+\frac{\nu}{2}\|u-u_{h}^{k}\|_{V}^{2}\\ \leq\langle E^{\prime}(u_{h}^{k}),\Pi_{h}u-u\rangle.$
@@
Proceeding, we multiply by $\alpha^{k}$, sum from $k=1$ to $k=\ell$, and use the energy dissipation property Lemma 3.2. This yields
(63) $(E(u_{h}^{\ell})-E(u))\sum_{k=1}^{\ell}\alpha_{k}+\frac{\nu}{2}\sum_{k=1}^{\ell}\alpha_{k}\|u-u_{h}^{k}\|_{V}^{2}+(\mathcal{D}(Bu,o_{h}^{\ell}),1)_{\Omega_{d}}\\ \leq(\mathcal{D}(Bu,o_{h}^{0}),1)_{\Omega_{d}}+\sum_{k=1}^{\ell}\alpha_{k}\langle E^{\prime}(u_{h}^{k}),\Pi_{h}u-u\rangle.$
We now define the weighted average $\overline{u}_{h}^{\ell}=\sum_{k=1}^{\ell}\alpha_{k}u_{h}^{k}/\sum_{k=1}^{\ell}\alpha_{k}$ and use Jensen’s inequality
(64) $\|u-\overline{u}_{h}^{\ell}\|_{V}^{2}\leq\frac{\sum_{k=1}^{\ell}\alpha_{k}\|u-u_{h}^{k}\|_{V}^{2}}{\sum_{k=1}^{\ell}\alpha_{k}}.$
@@
Thus, upon dividing 63 by $\sum_{k=1}^{\ell}\alpha_{k}$ then adding and subtracting $\langle E^{\prime}(u),\Pi_{h}u-u\rangle$, we obtain